OpenStudy (anonymous):

Is x= -1.75 the answer to this? I have already worked it out. I am just checking it. 4x+2(x+1)=9

hartnn (hartnn):

6x +2 = 9 6x = 7 x= 7/6 x= 1.1667 how did u get -1.75 ?

OpenStudy (anonymous):

first plug in your answer and if the answer matches up to 9 then thats your answer

OpenStudy (anonymous):

4x+2(x+1)=9 4x+2-2(x+1)=9-2 4x(x+1)=7 4x=7 4x/4=7/4 x=1.75 Sorry I meant to put positive 1.75 not negative. But I still have a feeling that my answer is incorrect.

OpenStudy (anonymous):

well take 1.75 and plug it in for x and what do you get

OpenStudy (anonymous):

not 9. so im incorrect

hartnn (hartnn):

4x+2(x+1)=9 4x+2-2(x+1)=9-2 itself is incorrect. why did u subtract 2 ? and its not done properly

OpenStudy (anonymous):

okay. so what is the first step to do? I have not had math in about a year so I need some refreshing. My teacher gave us practice problems for the first day, so she didn't teach or go over anything, so I am very lost.

hartnn (hartnn):

4x+2(x+1)=9 1st step is to simplify 2(x+1) by distributing